OpenStudy (anonymous):

Simplify the expression using the order of operations. 5+4^2/2(3)

OpenStudy (anonymous):

Is it \[\frac{ 5+4^2 }{ 2*3 }\] or \[5+\frac{ 4^2 }{ 2*3 }\]

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

Okay, so first, deal with the exponent. 4^2 would be 4*4, which is 16. So that leaves the top of the equation to be 5+16, which is equal to 21.

OpenStudy (anonymous):

The bottom of the equation is just a simple multiplication of 2*3, so that's six. So the entire thing would be 21/6.

OpenStudy (anonymous):

The answer is apparently 29 but I don't know why, I also got 21/6=3.5

OpenStudy (anonymous):

29? Hm... Well... Hang on, give me a moment to work this a different way.

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

I've tried a couple of other ways, but everything I'm getting isn't even near the 20's range. Sorry. =/

OpenStudy (anonymous):

Its okay :) thanks for your help.